# SAM2-ELNet: Label Enhancement and Automatic Annotation for Remote Sensing Segmentation

### v1

@@ III-A Encoder, Adapter, and RFB Modules
III-B Decoder
The original decoder follows the classical U-Net design, consisting of three decoder blocks, each comprising two Conv-BN-ReLU layers. When processing remote sensing images with blurred boundaries or low contrast, this structure tends to produce segmentation outputs with unclear contours, fragmented edges, or excessive smoothing. To address these issues, we introduce the edge attention mechanism (EAM) to enhance segmentation precision, particularly in boundary regions.
Figure 2: Architecture of the Hiera. Proposed in SAM2-Unet, the encoder leverages the Hiera backbone for multi-scale feature modeling, with adapters enhancing feature adaptation.
EAM enhances the feature map’s focus on boundary regions during the upsampling stage. After each upsampling step, it ensures that multi-scale feature aggregation preserves critical boundary information. The EAM in this study employs convolution operations to extract object boundary features and applies element-wise weighting to refine the input features, guiding the model to emphasize boundary details. The edge attention weight is computed as follows:
@@ III-B Decoder
where $\odot$ represents element-wise multiplication. $X^{\prime}$ is an enhanced feature map emphasizing edge-related information. This module adopts a lightweight design, using only a independent 3×3 convolutional layer, which significantly reduces computational cost compared to self-attention mechanisms.
III-C Loss Function
@@
We formulate the loss function as a combination of the weighted Intersection over Union (IoU) loss $L_{\text{wIoU}}$ and the weighted binary cross-entropy (BCE) loss $L_{\text{wBCE}}$, defined as:
$L=L_{\text{wIoU}}+L_{\text{wBCE}}$ (3)
@@
To enhance training effectiveness, we apply deep supervision to all segmentation outputs $S_{i}$, leading to the final total loss function:
$L_{\text{total}}=\sum_{i=1}^{3}L(G,S_{i})$ (4)
where $G$ denotes the ground truth labels, and $S_{i}$ represents the multi-level segmentation outputs generated by the model.
Figure 3: Labeling Module Architecture. Converts grayscale predictions into binarized labels using edge detection, morphological processing, and the watershed algorithm for precise region extraction.
III-D Labeling Module
The model’s initial predictions are grayscale images with intensity values ranging from 0 to 255. To transform these predictions into binarized labels, our work design a labeling module based on edge detection and the watershed transform. This approach integrates Canny edge detection [41], morphological processing [42], [43], distance transform, and the watershed algorithm [44], [45] to accurately extract foreground, background, and boundary regions. The structure of the labeling process is illustrated in Fig. 3.
To detail the labeling procedure, we apply the Canny operator to an input image $I$ to extract object contours, emphasizing prominent edges. This process produces the edge map $E$, defined as:
$E=\text{Canny}(I,T_{\min},T_{\max})$ (5)
where $E$ represents the binary edge map, while $T_{\min}$ and $T_{\max}$ denote the lower and upper threshold values of the Canny operator. This step enhances object boundaries by highlighting regions with sharp intensity transitions.
To fill gaps within foreground objects, we apply a combination of binarization and morphological dilation. After thresholding the input image, we perform dilation:
$T(x,y)=\begin{cases}255,&I(x,y)<T_{\text{thresh}}\\ 0,&\text{otherwise}\end{cases}$ (6)
$D=\text{dilate}(T,K,n)$ (7)
where $T_{\text{thresh}}$ is the binarization threshold, $K$ is the morphological kernel, and $n$ represents the number of dilation iterations. This operation enhances the foreground region, refines boundary continuity, and reduces discontinuities within the object.
To generate the segmentation labels, we construct the foreground mask $F_{\text{fg}}$ by merging the dilation result $D$ with the Canny edge map $E$, Similarly, we obtain the background mask $F_{\text{bg}}$ by inverting the binarization result:
$F_{\text{fg}}=D\cup E,\quad F_{\text{bg}}=\neg T$ (8)
Next, we optimize the foreground segmentation by constraining it with the Euclidean distance transform:
$D_{T}=\text{distanceTransform}(F_{\text{fg}},L_{2},5)$ (9)
$F_{\text{fg}}^{\prime}=\begin{cases}255,&D_{T}(x,y)>\alpha\cdot\max(D_{T})\\ 0,&\text{otherwise}\end{cases}$ (10)
where $\alpha$ is the threshold scaling factor, which controls the extent of the foreground region to ensure boundary precision without excessive expansion.
Using the extracted foreground and background regions, the labeling matrix $M(x,y)$ is constructed as:
$M(x,y)=\begin{cases}2,&(x,y)\in F_{\text{fg}}^{\prime}\\ 1,&(x,y)\in F_{\text{bg}}\end{cases}$ (11)
This matrix serves as the marker input for the watershed transform, which segments the image into foreground, background, and boundary regions based on gradient-based region growing:
$M^{\prime}=\text{Watershed}(I,M)$ (12)
The final labeled output $M^{\prime}$ assigns 0 for boundary regions and background, sets 255 for foreground objects, thereby generating high-quality segmentation labels.
III-E Overview of the Label Enhancement Process
@@
The overall workflow of the proposed label enhancement method consists of the following stages:
@@
    Initial Manual Annotation. A small subset of manually annotated data serves as the initial training set. Although these annotations may exhibit suboptimal quality, they still provide essential supervisory signals. This dataset supports the preliminary training of the model, enabling it to learn fundamental segmentation patterns and adapt to the specific characteristics of the target task.
@@
    Model Fine-Tuning. The model initializes with pretrained weights and fine-tunes on the manually annotated dataset. With the encoder weights remaining frozen, the fine-tuning process primarily optimizes the adapters and decoder. This strategy mitigates the adverse impact of suboptimal quality annotations on model learning, facilitating better adaptation to remote sensing segmentation tasks.
@@
    Label Generation. Upon completion of training, the fine-tuned model infers segmentation results, automatically generating high-quality segmentation labels.
@@
    Iterative Optimization (Optional). The model can further fine-tune using the enhanced labels, iteratively refining segmentation quality through adaptive learning. This iterative strategy progressively improves annotation accuracy and stability.
@@ IV-D1 Qualitative Analysis
The model fine-tunes using only the training subset of the PALSAR dataset for 200 epochs. After training, the model enhances both the training and test set labels of the PALSAR dataset, generating high-quality annotations for subsequent analysis.
From a qualitative perspective, the enhanced labels show substantial visual improvements over the original annotations, as illustrated in Fig. 5.
@@
(b) manual annotation
(c) enhanced label
@@
(e) manual annotation
(f) enhanced label
Figure 5: Results of the label enhancement experiment. Each triplet consists of an original SAR image, a manually annotated label, and a model-enhanced label. The enhanced labels exhibit smoother and more continuous boundaries, improved annotation details, and refined structural consistency.
As depicted in Fig. 5, the enhanced labels exhibit smoother and more continuous boundaries, significantly refining annotation details compared to the original labels. This improvement is not limited to isolated cases but is a consistent trend across the dataset. The enhanced labels demonstrate superior performance in boundary smoothness, object shape precision, and the accurate depiction of complex regions and ambiguous edges, enabling a more precise delineation of target contours. The increased consistency and accuracy of the annotations validate the model’s ability to effectively optimize manually labeled data, offering a more reliable supervisory dataset for remote sensing image segmentation tasks.
IV-D2 Quantitative Analysis
@@
In this experiment, we conduct comparative studies using the open-source implementation of DeepLabv3+ [46], [47] to analyze the impact of annotation quality on model performance. Specifically, we compare the performance of models trained on original labels versus those trained on enhanced labels. All experiments follow identical training conditions, utilizing DeepLabv3+ MobileNet, DeepLabv3+ ResNet50, and DeepLabv3+ ResNet101 as baseline models. Each model is initialized with the same pretrained weights and trained with a consistent set of hyperparameters for 20 epochs.
TABLE I: Impact of Different Label Qualities on Model Performance (DeepLabv3+ Training Comparison)
Model Label Type ACC $\uparrow$ mIoU $\uparrow$ Final Loss $\downarrow$
MobileNet Original Label 0.9251 0.7436 0.0773
Enhanced Label 0.9523 0.8164 0.0452
ResNet50 Original Label 0.9195 0.7396 0.0815
Enhanced Label 0.9476 0.8092 0.0458
ResNet101 Original Label 0.9210 0.7652 0.0792
Enhanced Label 0.9337 0.8136 0.0453
Experimental results Table I and Fig. 6 indicate that models trained with enhanced labels consistently outperform those trained with original labels in terms of mean Intersection over Union (mIoU) and overall accuracy (ACC). This demonstrates that enhanced labels facilitate more effective learning of representative features. Furthermore, models trained with enhanced labels converge faster and achieve lower final loss values, confirming that enhanced labels more accurately capture the true data distribution. As a result, leveraging enhanced labels significantly enhances overall performance in segmentation tasks.
(a) overall loss curve
(b) detailed loss segment
(c) overall accuracy curve
(d) overall mIoU curve
(e) initial loss curve
Figure 6: Training Convergence and Performance Metrics of DeepLabV3+ ResNet50 on Original and Enhanced Labels. (Loss, accuracy (ACC), and mean Intersection over Union (mIoU) trends during training on original and enhanced labels. Each epoch consists of 190 iterations, resulting in a total of 3,800 iterations over 20 epochs.)
IV-E Generalization Performance Evaluation
To evaluate the model’s generalization ability, we fine-tuned it solely on the PALSAR dataset’s training set, optimizing it over 200 epochs. After training, we used the fine-tuned model to refine both the training and testing labels of the Sentinel dataset, producing higher-quality annotations. This approach helped us assess the model’s adaptability to different data distributions and explore its potential for automatic annotation tasks.
@@ IV-E1 Qualitative Analysis
To visually assess the impact of label enhancement on segmentation performance in the Sentinel dataset, we conducted a comparative analysis through visual inspection. As shown in Fig. 7, models trained with enhanced labels outperform those trained on original labels, demonstrating sharper boundaries, finer segmentation details, and better preservation of target regions. The enhancement process effectively mitigates issues like boundary ambiguity and object omission, resulting in more precise segmentation outcomes.
(a) original image
(b) manual annotation
(c) enhanced label
(d) original image
(e) manual annotation
(f) enhanced label
Figure 7: Generalization Performance of Label Enhancement on the Sentinel Dataset. The enhanced labels improve boundary sharpness and segmentation accuracy, demonstrating the model’s adaptability to different data distributions.

### v2

@@ III-A Label Generation Network
In remote sensing image segmentation tasks, accurately delineating object boundaries plays a critical role, especially when the target objects such as roads, marine oil spills, or man-made structures appear narrow, elongated, or fragmented. Inaccurate boundary handling can lead to serious consequences. For instance, boundary errors in road segmentation may disrupt connectivity analysis, affecting downstream applications like navigation or path planning. In oil spill detection, boundaries often lack clarity and exhibit smooth intensity transitions, making them hard to distinguish from the background. Without adequate boundary awareness, models tend to produce overly expanded prediction regions, significantly overestimating the affected area and misleading environmental assessments or emergency responses.
The baseline decoder follows the classical U-Net design, consisting of three decoding modules, each with two Conv-BN-ReLU layers. However, this structure struggles with blurred boundaries or low-contrast regions. To address this issue and enhance the model’s ability to capture fine boundary details, we introduce the EAM. Unlike traditional convolutional modules that aggregate spatial features uniformly, EAM focuses on the transition zones between foreground and background. It strengthens feature responses in edge regions and improves boundary awareness. The network applies EAM after each upsampling step to ensure that multi-scale feature fusion retains essential boundary cues. Specifically, EAM uses convolutional layers to extract edge-related features and applies element-wise weighting to refine the input features, guiding the model to emphasize edge information. The attention weights for edge refinement are computed as follows:
@@
where $\odot$ represents element-wise multiplication. $X^{\prime}$ is an enhanced feature map emphasizing edge-related information. This module adopts a lightweight design, using only an independent 3×3 convolutional layer, which significantly reduces computational cost compared to self-attention mechanisms.
III-B Loss Function
@@
We formulate the loss function as a combination of the weighted Intersection over Union (IoU) loss $L_{\text{wIoU}}$ and the weighted binary cross-entropy (BCE) loss $L_{\text{wBCE}}$, controlled by a balancing factor $\lambda$, defined as:
$L=\lambda\cdot L_{\text{wIoU}}+(1-\lambda)\cdot L_{\text{wBCE}}$ (3)
@@
To enhance training effectiveness, we apply deep supervision to all segmentation outputs $S_{i}$, assigning each output a weight $\alpha_{i}$. The final total loss function is defined as:
$L_{\text{total}}=\sum_{i=1}^{3}\alpha_{i}\cdot L(G,S_{i})$ (4)
where $L(G,S_{i})$ denotes the loss between the ground truth $G$ and the predicted segmentation output $S_{i}$, and $\alpha_{i}$ is the supervision weight assigned to each level $i$ of the segmentation output. The weights $\alpha_{i}$ are normalized such that $\sum\alpha_{i}=1$.
III-C Label Quality Evaluator
Automatically generated labels inevitably contain noise and bias in downstream applications. To ensure label reliability, this study introduces a label quality evaluator that assesses the consistency of predicted labels and filters out unstable or low-confidence results. The structure of the evaluator is illustrated in Fig. 4.
Each remote sensing image $I$ undergoes three types of basic data augmentation, such as Gaussian noise, horizontal flipping, and random rotation, resulting in three perturbed versions denoted as $I_{1},I_{2},I_{3}$. A label generation network with different fine-tuning epochs processes the three augmented inputs separately, producing segmentation predictions $P_{1},P_{2},P_{3}$. The three outputs undergo a mutual alignment process to ensure that all predictions share the same spatial orientation, enabling consistent pixel-wise comparison.
To quantify pixel-level similarity among the binarized predictions, the root mean squared error (RMSE) is used as the consistency metric. The RMSE score $R_{i}$ at each pixel location $i$ is calculated as:
$R_{i}=\sqrt{\frac{1}{3}\sum_{j=1}^{3}\left(p_{i}^{(j)}-\bar{p}_{i}\right)^{2}}$ (5)
In this formulation, $p_{i}^{(j)}\in\{0,1\}$ denotes the binary foreground prediction at pixel $i$ in the $j$-th perturbed input. The mean prediction $\bar{p}_{i}$ is computed as:
$\bar{p}_{i}=\frac{1}{3}\sum_{j=1}^{3}p_{i}^{(j)}$ (6)
The RMSE reflects the level of disagreement at pixel $i$. A lower $R_{i}$ indicates stronger agreement among the three predictions, suggesting a more reliable label. To assess the overall consistency of a label, the final RMSE score is obtained by averaging $R_{i}$ over all pixel positions.
Global consistency across binary segmentation masks is measured using the Intersection over Union (IoU) and Dice coefficient. For any two masks $A$ and $B$, these metrics are computed as:
$\text{IoU}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (7)
$\text{Dice}(A,B)=\frac{2|A\cap B|}{|A|+|B|}$ (8)
Average IoU and Dice scores across all three pairs of predictions $P_{1},P_{2},P_{3}$ are calculated using the following expressions:
$\text{IoU}_{\text{avg}}=\frac{1}{3}\sum_{1\leq i<j\leq 3}\text{IoU}(P_{i},P_{j})$ (9)
$\text{Dice}_{\text{avg}}=\frac{1}{3}\sum_{1\leq i<j\leq 3}\text{Dice}(P_{i},P_{j})$ (10)
A label is considered low-quality if the pixel consistency score $C_{i}$, the average IoU, or the Dice score remains below a predefined threshold.
To obtain an overall label quality score, the pixel-wise RMSE score $R$, average IoU, and average Dice coefficient are combined using a weighted formulation:
$Q=\beta_{1}\cdot(1-R)+\beta_{2}\cdot\text{IoU}_{\text{avg}}+\beta_{3}\cdot\text{Dice}_{\text{avg}}$ (11)
The weighting coefficients $\beta_{1},\beta_{2},\beta_{3}$ reflect the relative contribution of each metric. The RMSE term is transformed into 1R1-R to ensure that a higher score consistently indicates better agreement among the predictions.
A label is marked as low-quality and scheduled for re-annotation if the final quality score $Q$ falls below a predefined threshold.
III-D Overview of the Label Enhancement and Automatic Annotation Process
@@
The proposed label enhancement and automatic annotation process consists of the following stages:
@@
    Initial Label Preparation. Prepare a small set of high-quality manually annotated samples when performing automatic annotation, or construct a large set of coarse manual annotations when conducting label enhancement. These datasets serve as the initial supervision sources for model fine-tuning.
@@
    Model Fine-Tuning. Fine-tune the model using the prepared initial labels. The model initializes with pretrained weights, while only the adapters and decoder are updated during training to preserve the generalization ability of the encoder.
@@
    Automatic Label Generation and Filtering. Apply the fine-tuned model to generate segmentation labels. The label quality evaluator filters out low-quality predictions, which are flagged for re-annotation to ensure label reliability.
@@
    Model Refinement with Filtered Labels. Incorporate the retained high-quality labels into the training set and further fine-tune the model. This refined model can generate more accurate annotations in subsequent iterations.
  5. 5.
    Iterative Enhancement. Repeat the label generation, quality filtering, and model refinement steps iteratively. This progressive process incrementally improves annotation quality and model performance in resource-constrained scenarios.
@@ IV-A1 Deep-SAR Oil Spill (SOS) dataset
IV-A2 CHN6-CUG Road dataset
The CHN6-CUG Road dataset [56] is a high-resolution optical road segmentation dataset constructed from Google Earth imagery. It covers six representative cities in China with varying urban characteristics: Beijing (Chaoyang), Shanghai (Yangpu), Wuhan, Shenzhen (Nanshan), Hong Kong (Sha Tin), and Macao. The dataset includes 4,511 manually annotated images with a resolution of 50 cm/pixel and size $512\times 512$, of which 3,608 are used for training and 903 for testing. Labeled roads encompass various transportation types, including railways, highways, urban, and rural roads. Due to occlusions and complex urban structures, the annotations may exhibit challenges such as partial coverage and blurred boundaries, making this dataset suitable for evaluating segmentation robustness under realistic conditions.
@@ IV-D1 Qualitative Analysis
To simulate the practical application scenario of “low-cost coarse annotation + label enhancement” we treat the label annotations in the Deep-SAR Oil Spill (SOS) dataset as coarse labels to be enhanced. Both quantitative and qualitative evaluations are conducted accordingly.
The model is fine-tuned using only the training subset of the PALSAR dataset for 200 epochs. After training, it enhances both the training and test set labels of the PALSAR dataset, generating high-quality annotations for subsequent analysis.
As shown in Fig. 6, the enhanced labels exhibit smoother and more continuous boundaries, and significantly refine annotation details compared to the original labels. This improvement appears consistently across the dataset. The enhanced annotations provide better object shape precision and more accurate representation of complex regions and ambiguous edges, enabling clearer delineation of target contours. These results demonstrate the model’s ability to effectively optimize manual annotations.
@@
(b) manual
(c) U-Net
(d) SAM2(Guidance)
(e) SAM2-UNet
(f) LOGCAN++
(g) Ours(enhanced)
($a$)
Figure 6: Comparison of label enhancement results across different methods. Each group shows the original SAR image, the manual annotation, and enhanced labels generated by our model, U-Net (same pipeline), SAM2 (guided by manual labels), SAM2-UNet, and LOGCAN++. Our method produces labels with finer detail, more accurate boundaries, and overall higher annotation quality, achieving comparable performance to current state-of-the-art approaches.
To further evaluate the model, we trained a U-Net using the same pipeline. U-Net [57] showed moderate improvements, but our model produced more precise and consistent annotations. We also evaluated SAM2 [25] in two variants: SAM2 (guided by manual labels) and SAM2-UNet. Compared to SAM2 (guided by manual labels), our model achieved noticeably better boundaries and structure. Compared to SAM2-UNet, it provided slightly finer details and more stable quality. In addition, we compared our approach with the current SOTA segmentation model LOGCAN++ [58], and achieved comparable performance, further demonstrating the effectiveness of our method in label quality enhancement tasks.
IV-D2 Quantitative Analysis
Our objective is to generate higher-quality annotations. Directly evaluating models using only the original labels may fail to fully reflect the effectiveness of the enhanced annotations. Moreover, since segmentation models tend to exhibit low sensitivity to label quality during training, the presence of noisy annotations may even improve model robustness. To address these limitations, we design a new evaluation strategy that constructs multiple test sets to more accurately assess annotation quality.
TABLE I: Label Quality Evaluation Using Third-Party Segmentation Models
Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
Test-Orig 0.7541 0.0274 0.9157 0.0127
Test-Enh 0.7877 0.0062 0.9328 0.0045
U-Net Original Test-HQ 0.8276 - 0.9462 -
Test-Orig 0.7919 0.0357 0.9350 0.0112
Test-Enh 0.8369 0.0093 0.9491 0.0029
TABLE II: Evaluation of Enhanced Labels Using Baseline Models
Test Model Train Set Test Set mIoU $\Delta$ from HQ (mIoU) ACC $\Delta$ from HQ (ACC)
DeepLabv3+ Original Test-HQ 0.7815 - 0.9283 -
Test-U-Net 0.8517 0.0701 0.9466 0.0182
Test-SAM2-UNet 0.8192 0.0377 0.9360 0.0076
Test-LOGCAN++ 0.7893 0.0077 0.9324 0.0041
Test-Ours 0.7877 0.0062 0.9328 0.0045
Specifically, a group of domain experts were invited to refine the annotations of selected samples in the test set, forming a high-quality benchmark test set referred to as Test-HQ. Corresponding samples from the original and enhanced labels were also selected to construct two comparison test sets, Test-Orig and Test-Enh. A third-party segmentation model undergoes training with the original training set and evaluation on the three test sets. Performance metrics offer an indirect means to evaluate the actual quality of both original and enhanced labels. Although traditional metrics may not fully reflect the quality of annotations, the performance gap between these results and Test-HQ enables a more objective assessment of label quality.
(a) Loss
(b) mIoU
(c) Accuracy
Figure 7: Evaluation results of the model trained on original labels across three different test sets. The performance on Test-Enh is notably closer to that on Test-HQ, indicating the higher quality of the enhanced annotations.
(a) original image
(b) manual labels
(c) auto-labels
@@ IV-D2 Quantitative Analysis
(e) manual labels
(f) auto-labels
Figure 8: Generalization performance of automatic annotation on the Sentinel dataset. The automatically generated labels not only preserve key target structures but also exhibit improved quality in certain regions compared to the original annotations.
@@
In this experiment, we adopt open-source implementations of DeepLabv3+ and U-Net [55, 59, 57] to analyze the quality of the generated annotations. Specifically, the performance of models trained on original labels is evaluated across the three test sets, and the performance gap relative to Test-HQ is calculated to assess label quality more accurately. All experiments are conducted under the same training conditions. The backbone networks for DeepLabv3+ include MobileNet, ResNet50, and ResNet101. Each model is initialized with the same pretrained weights and trained for 20 epochs using consistent hyperparameters.
The experimental results in Table I and Fig. 7 show that models trained on the original labels perform significantly better on both Test-Enh and Test-HQ in terms of mean Intersection over Union (mIoU) and accuracy (ACC), compared to their performance on Test-Orig. Moreover, the results on Test-Enh are closely aligned with those on Test-HQ, further demonstrating that the enhanced labels have achieved a quality level comparable to expert-level fine annotations.
To verify the effectiveness of our method, we applied the same evaluation protocol to several representative baseline models. The results are summarized in Table II.
When using model-generated labels as test sets, evaluation scores often exceed those obtained with manual annotations. This happens because coarse or overly smooth test labels make it easier for the model to produce matching predictions, thus inflating the evaluation metrics. Visual inspection reveals that labels generated by U-Net tend to cover broader regions. When used as test labels, they yield higher scores and exhibit a larger gap from the high-quality reference set (Test-HQ).
In contrast, our method consistently delivers results that are closer to Test-HQ, indicating that the generated labels better approximate expert-level fine annotations. Furthermore, our approach achieves comparable performance to the state-of-the-art model LOGCAN++, reinforcing the reliability of the proposed evaluation scheme and confirming the effectiveness of the label enhancement framework.
(a) Loss
(b) mIoU
(c) Accuracy
Figure 9: Comparison of model performance trained on auto-annotated versus original labels. The results demonstrate comparable performance, confirming the practical effectiveness of automatic annotation.
(a) original image
(b) manual labels
(c) auto-labels
(d) original image
(e) manual labels
(f) auto-labels
Figure 10: Visual Comparison of Auto-Annotated and Manually Labeled Data. The auto-annotated labels exhibit comparable quality to manual annotations, effectively capturing key structures and fine boundaries in high-resolution remote sensing images.
IV-E Automatic Annotation Performance Evaluation
To further evaluate the generalization capability of the proposed model, particularly its potential in automatic annotation tasks, we design a new set of experiments to simulate a typical cross dataset annotation scenario. Specifically, the model is fine-tuned on the training set of the PALSAR dataset for 200 epochs, and then used to automatically generate labels for the training set of the Sentinel dataset.
@@ IV-E1 Qualitative Analysis
To visually assess the quality of automatically generated labels in the Sentinel dataset, we conducted a comparative analysis through visual inspection. As shown in Fig. 8, the automatically generated labels demonstrate superior quality compared to the original annotations, characterized by clearer boundaries, finer details, and improved representation of target regions.